OpenStudy (anonymous):

Lee needed to make drapes the length from the floor to the top of the window. The wall beneath the window had the height of 2 2/3 feet and the height of just the window measured 2 3/4 feet. How long do the drapes have to be?

OpenStudy (johnweldon1993):

So we need to add the 2 lengths together \[\large 2\frac{2}{3} + 2\frac{3}{4}\] First change to improper fractions \[\large \frac{8}{3} + \frac{11}{4}\] Now find a common denominator \[\large \frac{32}{12} + \frac{33}{12}\] Now just add over the common denominator \[\large \frac{32 + 33}{12} = \frac{65}{12}\] And finally lets just change that back to a mixed number \[\large 5\frac{5}{12}\]